Two cars leave an intersection at the same time. One is headed south at a constant speed of 30 miles per hour; the other is headed west at a constant speed of 40 miles per hour. What is the distance between the cars in 30 minutes?
Answer by mananth(16946)

southbound speed = 30 mph
westbound speed = 40 mph
...
in half hour they have traveled
15 miles and 20 miles repectively.
Their position and the intersection form a right triangle
..
15^2+20^= D^2
225+400=D^2
625= D^2
D= sqrt 625
D= 25 miles